ESTIMATING THE ASYMPTOTIC DISPERSION OF THE L_{1} MEDIAN

ARUP BOSE

Theoretical Statistics and Mathematics Unit, Indian Statistical Institute, 203, B.T. Road, Calcutta-700 035, India

(Received February 21, 1994; revised July 27, 1994)

Abstract. A simple estimate of the asymptotic dispersion matrix of the L_{1} median is proposed and its rate of convergence is studied.

Key words and phrases: L_{1} median, asymptotic dispersion, consistent estimate, rate of convergence, strong law of large numbers, central limit theorem, law of iterated logarithm.